6 % of the 100 students in Emmie Xi Hem International School students said that they had been late in their first-period classes. What is the upper bound of the 99% confidence interval for the proportion of students in the said school who had been late in their first-period classes?


1
Expert's answer
2022-03-28T16:05:18-0400

"\mu +z\sqrt{\frac{p(1-p)}{n}}=0.06+2,5758 \sqrt{\frac{0.06(1-0.06)}{100}}=0.12"